# EpidemIQs: Prompt-to-Paper LLM Agents for Epidemic Modeling and Analysis

### v1

@@ V-A Review Score and Computational Cost
TABLE II: Papers Evaluation of Five Epidemic Questions by AI and Human Reviewers
Question Topics
Readability
Relevance
& Motivation
Technical
Soundness
Experimental
Rigor
Limitation
& Ethics
Overall
AI Human AI Human AI Human AI Human AI Human AI Human
1. Effect of degree heterogeneity in SEIR model 9.5 7 9.5 9 9.5 8 10 8 9.5 8 9.5 8.0
2. Reasons of transmission break 9.5 7 10 10 10 6 9 10 8 9 9.7 8.2
3. Effect of temporality on epidemic behavior in activity-driven networks 9.5 9 9.5 9 9 7 9.5 7 8 7 9.2 7.8
4. Coexistence of exclusive bi-virus spread on multiplex networks 9 7 9 9 9.5 9 9.3 8 8.2 7 9.0 8.0
5. Herd-immunity through random and targeted vaccination in networks 9 6.5 9 9 10 8 10 9 8 8 9.2 8.1
Table II presents the evaluation scores assigned by an LLM-as-Judge and by human experts across five rubric criteria. For each of the five epidemic questions, the framework autonomously generated ten full papers, yielding 50 manuscripts. We randomly selected five of these for blind human review. Each reviewer scored every criterion on a 1‑to‑10 scale. Human scores averaged 7.98 with a standard deviation of 0.20 ($\approx 2.5\%$), indicating consistently strong performance by EpidemIQs. Reviewers praised the manuscripts’ comprehensive problem framing, robust experimental design and analysis, and balanced coverage of both analytical and stochastic aspects. The main criticisms concerned verbosity with occasional repetition of content, the use of uncommon vocabulary that slowed reading, and some missing or incomplete figure references.
Table III reports the average turnaround time, total number of tokens, and total cost (calculated by OpenAI’s May 2025 pricing) for all questions, measured from the initial query to the final report. A total cost of only \$1.57, a turnaround of under 30 minutes, a 100% success rate, and consistently high review scores demonstrate the strength of the framework and the efficiency of its multi‑agent orchestration. The low cost and rapid execution are achieved by splitting tasks between expert and scientist agents. Token‑heavy, low‑complexity jobs are delegated to fast, inexpensive models such as GPT‑4.1 mini, while more demanding works such as planning, reasoning, tool coordination, and output refinement is handled by the full GPT‑4.1 model.
TABLE III: Average Computation Statistics
@@
V-B Multi-Turn Single-Agent vs EpidemIQs Orchestration
In this section, we ran the single-agent across the same five questions ten times, using two OpenAI LLMs, GPT‑4.1 and o3. Single-agent‑GPT‑4.1 achieved an average completion success rate of 78±7.7% and an average human-review score of 5.06, whereas single-agent-o3 reached 80.0±6.32% with a human-review score of 5.68. The LLM-as-Judge provides comparative evaluations of the successful papers generated by single-agent and EpidemIQs. Table IV presents these scores for all questions.
TABLE IV: Overall Score for AI and Human Expert Evaluation of Generated Papers
@@ V-B Multi-Turn Single-Agent vs EpidemIQs Orchestration
EpidemIQs 9.5 8.0 9.5 8.2 9.2 7.8 9.0 8.0 9.2 8.1
single-agent-GPT4.1 8.9 5.4 7.3 5.3 8.9 4.8 7.5 4.6 8.4 5.2
single-agent-o3 8.5 5.8 8.6 6.3 8.3 5.5 8.6 5.8 7.4 5.2
Surprisingly, the AI-generated scores differ only slightly across the three cases, while the human reviews reveal substantial disparities in their outcomes. The single-agent-GPT‑4.1 setup addressed the first three questions to some extent; however, it was limited: it handled the analytical component inadequately, ran only a few stochastic simulations, and in some instances merely solved the ODE system instead of performing exact stochastic simulations. For the more complex question 5, none of the trials identified the target vaccination threshold or model the network appropriately. Furthermore, the parameters and the rationale for their selection were not discussed comprehensively.
@@
The conclusions were sometimes misleading, as they were not based on the simulation results; instead, the agent often relied on prior knowledge or literature. When grounded in results, they were still occasionally inaccurate, for example, in the targeted-vaccination case, the agent typically claimed that vaccinating all high-degree nodes was necessary, largely because it generated the network before calculating the analytical thresholds, resulting in very sparse high-degree nodes. Human reviews also mention that while the agent could address the task to some degree, it is in shallow depth, missing figures, insufficient simulations, and limited analyses, and the overall failure to address the tasks comprehensively, compared to rigorous analysis and experimentation in the MA setup.
On the other hand, The single-agent-o3 did better job with analytical aspect of the question and better writing quality, however, the simulations still was very limited, and the agent shows more autonomy in how to accomplish tasks rather based on the expert knowledge provided as instructions, for example for stochastic simulation it did not always use the engine provided for it or ignoring literature and online searches. Though the single-agent-o3 generally generated higher quality reports, there was still some cases when there were inaccuracies, for example for question four, even the coexistence was not achieved through stochastic simulation, it was concluded that simulation results confirm that coexistence is possible when analytical conditions are met or the agent mentioned the simulation were performed by FastGEMF over 1000 realizations, however scripts in the repository showed custom engine for deterministic simulation was used to simulated SI instead of SIR, which showed a huge drift from the main task and hallucination of the agent in the report. Also, in none of the cases could it correctly address the targeted vaccination in question five. Finally, while reviewers praised the writing quality, conciseness, and analytical aspects of the reports, they found that they fell short in having enough depth, simulation rigor, and failure to address complex tasks comprehensively.
@@
It is also necessary to mention that, to avoid violating OpenAI’s usage policy when using the o3 model, we changed the questions from virus spread to meme spread. Otherwise, the success rate of single-agent-o3 was less than 30%. We did not consider this API error in our benchmarks for evaluating the o3 success rate. It is important to emphasize that this change in context does not affect the validity of the comparative results, as epidemic spread can easily be generalized for social cases such when a meme spread over a population.
Overall, the human reviews and performance analysis demonstrated the superior performance of EpidemIQs compared to the single-agent implementation. Single-agent systems consistently failed to address complex scenarios, exhibited drift in their outcomes, lacked depth in their reports, and had a significantly lower success rate in completing tasks. Table V presents the computational statistics for these three different implementations. These comparisons are particularly meaningful given that the EpidemIQs achieved its superiority at a cost comparable to single-agent-GPT4.1 and at less than half the cost of single-agent-o3, despite the o3 model having greater reasoning and coding capabilities than those employed in the EpidemIQs, as mentioned in Table I. Moreover, the superiority of the proposed framework extends beyond performance and computational efficiency, offering greater interpretability by providing explicit reasoning for each step.
@@
TABLE V: Comparison of Average Computation Statistics Across All Questions Over Three Different Implementations
@@
V-C Discussion, Limitations, and Ethical Concerns
@@
Although the EpidemIQs framework consistently produces high‑quality manuscripts throughout the workflow, it is not designed to replace human authorship in scientific research, but rather to be an assistant to human experts to facilitate the research process for testing and implementing ideas in the epidemic modeling field. EpidemIQs can still make mistakes, and AI‑generated prose can be highly persuasive, as [57, 58, 59] discussed, so human oversight remains essential to safeguard accuracy and integrity, and generated papers are meant to provide a comprehensive and unified result of the framework, making it easy to analyze the outcomes.
@@
Across all of our experiments, we found no major instances of hallucination in EpidemIQs outputs, and performance was remarkably consistent. Since LLMs are non‑deterministic, they may tackle the same question slightly differently each time; however, no substantial drift or incorrect reasoning was observed. For example, in Question 5, ModelScientist sometimes designs an SIRV (V represents vaccinated state) model and at other times an SIR model, or NetworkScientist may construct networks with different degree distributions yet the same target centrality. These variations are valid and do not diminish performance; they simply reflect the natural diversity of approaches to a single problem. However, there were errors observed during the Analysis phase. For example, in question three, where DataExpert misinterpreted an average recovery count below one as a fraction rather than a discrete number of individuals, DataScientist detected this inconsistency and deferred to VisionExpert, whose interpretation matched both the simulation protocol and analytical expectations. While such checks limit the impact of isolated calculation errors, occasional mistakes by DataExpert indicate that automated data analysis may be a part of further improvement.
@@
The framework also demonstrated sound awareness of its computational constraints. For example, in question three, we explicitly restricted simulations to the FastGEMF engine. SimulationScientist adhered to this limitation, employing FastGEMF exclusively for static networks, and it was explicitly noted in the final report that software limitations prevented temporal network experiments. This transparency is critical, as it prevents the framework from falsely claiming task completion. In contrast, such issues were observed in the single-agent implementation, where it falsely hallucinates task accomplishment. Furthermore, the reliability of scientific outcomes often hinges on the algorithms employed, which can be inherently complex. For instance, in one of the trails within question three, SimulationScientist designed a custom engine for simulating disease spread over a temporal network but erred in updating the order of states. This led to the premature recovery of initially infected nodes, preventing further spread and resulting in an underestimation of the final outbreak size and probability. Although DataScientist noted a significant discrepancy between simulated and expected analytical results, it incorrectly attributed the difference to the effects of temporality rather than the engine’s design flaw.
@@
To further investigate the agents’ attention to scientific details, in question three we explicitly specified the constant activity rate (and not activity probability per unit time), $\alpha=3$, which is unlikely in real-world case studies and rare in literature but was chosen to test the agents’ attention to subtle details. However, NetworkScientist creates the network edges with the probability of a node being activated as $p=\alpha\Delta t$, with $\Delta t=1$, as the first order Taylor series approximation of $p=1-\exp^{-\alpha\Delta t}$, where is only is acceptable if $\alpha\Delta t\ll 1$ (However, if $\alpha$ is probability $\alpha\Delta t$ is exact), making it invalid for $\alpha\Delta t=3>1$ resulting all nodes being activated each time, instead of having exact probability $\approx 0.95$.
@@
1) the importance and impact of providing appropriate tools to obtain consistent and reliable results, and 2) the continued need for a human expert to validate results, because agents can still make mistakes that are difficult to detect and may produce persuasive responses that justify incorrect conclusions.

### v2

@@ V-A Review Score and Computational Cost
Figure 8: Evaluation results of five epidemic questions by AI and human reviewers across different rubrics.
Figure 8 presents the evaluation scores assigned by an LLM-as-Judge and by human experts across five rubric criteria. For each of the five epidemic questions, the framework autonomously executed the pipeline 20 times, yielding 100 project repositories, including reports and artifacts. We randomly selected five of them for blind human review. Each reviewer scored every criterion on a 1‑to‑10 scale. Human scores averaged 7.98±0.20, indicating consistently strong performance by EpidemIQs. Reviewers praised the comprehensive problem framing in manuscripts, robust experimental design and analysis, and balanced coverage of both analytical and stochastic aspects. The main criticisms concerned verbosity with occasional repetition of content, the use of uncommon vocabulary that slowed reading, and some missing or incomplete figure references.
TABLE II: Average Per-Phase and Total Computation Statistics
@@
Table II reports the average turnaround time, total number of tokens, and total cost (calculated by OpenAI’s May 2025 pricing) for all questions, measured from the initial query to the final report. A total cost of only \$1.57, a turnaround of under 30 minutes, a 100% workflow completion rate, and high review scores demonstrate the strength of the framework and the efficiency of its multi‑agent orchestration. The low cost and rapid execution are achieved by splitting tasks between expert and scientist agents. Token‑heavy, low‑complexity jobs are delegated to fast, inexpensive models such as GPT‑4.1 Mini, while more demanding tasks such as planning, reasoning, tool coordination, and output refinement are handled by the full GPT‑4.1 model. To approximate the human-equivalent effort for EpidemIQs’ average performance, we adopt a similar approach from task duration baselines from Model Evaluation and Threat Research (METR) [50]. Our estimate is approximately 38.4 hours for a complete execution cycle, from literature retrieval to final reporting. A detailed breakdown is provided in Supplementary Section LABEL:S-sec:METR
V-B Ablation Study
To better understand what drives the EpidemIQs’ performance, we conducted an ablation study isolating the contribution of each scientist module. Next, evaluating the iterative single-agent performance offers a meaningful baseline for better understanding the added value of EpidemIQs orchestration mechanism.
V-B1 Scientist Agent Modules Analysis
To systematically analyze how different parts of the scientist agent contribute to performance, we conducted an ablation study on the DSBench on data analysis tasks [51]. DSBench is a recent benchmark for advanced data science tasks that requires multimodal inputs such as figures, text, and tabular data. In all configurations, we keep the backbone model fixed to OpenAI GPT 4.1 Mini, and we only change the internal architecture of the scientist agent. We first consider a simple LLM-only setting, where the model receives the question and the data but no explicit tool use or control over its own reasoning. In this case, the average accuracy over the 48 DSBench challenges is 29.76%. Adding the ReAct style tool use and intermediate reasoning increases accuracy to 37.40%, which shows that even a relatively light agent structure already helps the model use the data in a more systematic way. Next, we enable the planning module. This configuration, with planning and ReAct but zero reflection step, reaches 44.49%. When reflections were introduced with one step but without planning (ReAct/Reflect, reflection steps=1), performance reached 46.60%. Combining all modules with one reflection step (Plan/ReAct/Reflect, reflection steps=1) achieved 47.70%. Finally, we increase the reflection depth to five steps in the full configuration with planning, ReAct, and reflection. This full scientist architecture achieves 51.71% accuracy, which is the best result across. The gain from one to five reflection steps is smaller than the earlier jumps, suggesting that the benefits of additional reflection begin to plateau, though they continue to yield improvements in accuracy. Figure 9 summarizes the benchmark results. Overall, results show a mostly monotonic improvement when we move from a plain LLM-only agent to a structured scientist agent with tools, planning, and deeper reflection, and it confirms that all three modules are helpful in practice for non-trivial data analysis tasks, with 73% improvement of accuracy compared to using only plain LLM (from 29.76% to 51.71%).
Figure 9: Accuracy of different configurations of scientist agent over 48 DSBench data science challenges, using OpenAI GPT 4.1-Mini as backbone LLM. The systematic increase in accuracy (from 29.76% to 51.71%) demonstrates that incorporating planning and reflection modules significantly enhances the agent’s performance compared to a standalone LLM.
V-B2 Iterative Single-Agent vs EpidemIQs Orchestration
We ran the iterative single-agent across the same five questions ten times, using two OpenAI LLMs, GPT‑4.1 and o3. Single-agent‑GPT‑4.1 achieved an average workflow completion rate of 78±7.7% and an average human-review score of five sampled successful results of 5.06±0.32, whereas single-agent-o3 reached 80.0±6.32% with a human-review score of 5.68±0.37. The LLM-as-Judge provides comparative evaluations of the successful papers generated by single-agent and EpidemIQs. Table III presents these scores for all questions. The task success rate was 22±6% and 26±9.75% for GPT 4.1 and o3, respectively, as backbone LLMs. Specifically, in none of the trials could they achieve the correct targeted vaccination threshold in Question 5, and only about 10% of the time the stochastic coexistence regime in Question 4 was correctly validated.
TABLE III: Overall Score for AI and Human Expert Evaluation of Generated Results
@@ V-B2 Iterative Single-Agent vs EpidemIQs Orchestration
EpidemIQs 9.0±0.1 8.0 9.0±028 8.2 9.2±0.21 7.8 9.0±0.15 8.0 8.96±0.23 8.1
single-agent-GPT4.1 8.4±0.63 5.4 8.4±0.52 5.3 7.9±0.9 4.8 7.3±1.8 4.6 7.9±1.4 5.2
single-agent-o3 8.2±0.63 5.8 8.6±0.55 6.3 8.5±0.2 5.5 8.4±0.36 5.8 8.4±0.45 5.2
Human reviewers observed large performance gaps across the three implementations, even though the AI-generated scores were close. In the single-agent GPT-4.1 setting, the framework could partially address the first three questions, but it often handled the analytical part weakly, ran only a few stochastic realizations, and sometimes replaced exact stochastic simulations with solving ODEs. For the more complex Question 5, it did not identify the vaccination threshold and did not model the network in a suitable way, and the parameter choices were not well justified .
@@
Reviewers also noted that some conclusions were misleading as they were not supported by the actual simulation results, and the agent relied on prior knowledge rather than evidence. Even when results were used, they could be incorrect. For example, in targeted vaccination, the agent often claimed that vaccinating all high-degree nodes is necessary, partly because it generated a sparse network prior to computing thresholds. Overall, the single-agent reports were judged as shallow, with missing figures, limited simulations, and insufficient analysis compared to the multi-agent setup.
The single-agent o3 produced better analytical discussions and more polished writing, but the simulations remained limited, and it showed more drift in execution, for example, not consistently using the provided stochastic engine or not using literature and online search when needed. We found clear inconsistencies in some cases, such as claiming that coexistence was confirmed by stochastic simulations when it was not, or that FastGEMF conducted multiple realizations, whereas the repository’s Python scripts indicated a deterministic custom run and even an incorrect model in parts. Similar to GPT-4.1, it still failed to correctly solve the targeted vaccination problem in Question 5, and reviewers concluded that depth and experimental rigor were insufficient for the harder tasks.
@@
It is also necessary to note that, to avoid violating OpenAI’s usage policy when using the o3 model, we changed the questions from virus spread to meme spread. Otherwise, the workflow completion rate of single-agent-o3 was less than 30%. We did not consider this API error in our benchmarks for evaluating the o3 workflow completion rate. It is important to emphasize that this change in context does not affect the validity of the comparative results, as epidemic spread can easily be generalized for social cases, such as when a meme spreads over a population.
Overall, human reviews and the task level outcomes show that EpidemIQs addresses the questions more reliably than either single-agent baseline. The single-agent systems were more likely to drift, miss key requirements, and fail in complex scenarios, whereas EpidemIQs maintained greater depth and end-to-end completeness. Table IV reports the computation statistics, showing that these gains are achieved at a cost comparable to single-agent GPT-4.1 and at less than half the cost of single-agent o3, while also improving interpretability by making the reasoning steps explicit.
@@
TABLE IV: Comparison of Average Computation Statistics Across All Questions Over Three Different Implementations
@@
V-C Discussion of Results
@@
Although EpidemIQs demonstrated its applicability and produced high‑quality results and manuscripts throughout our evaluation, it is not designed to replace human authorship in scientific research. Its main goal is to serve as an assistant to human experts, facilitating the research process for testing and implementing ideas in epidemic modeling. EpidemIQs can still make mistakes (%79 task success rate), and AI‑generated prose can be highly persuasive, as [52, 53, 54] discussed, so human oversight remains essential to safeguard accuracy and integrity, and generated papers are meant to provide a comprehensive and unified result of the framework, facilitating the analysis of outcomes.
@@
Across all of our experiments, we found no major instances of hallucination in EpidemIQs outputs, and performance was remarkably consistent. Since LLMs are non‑deterministic, they may tackle the same question slightly differently each time; however, no substantial drift or incorrect reasoning was observed. For example, in Question 5, Model Scientist sometimes designs an SIRV (V represents vaccinated state) model, and at other times an SIR model, or Network Scientist may construct networks with different degree distributions yet the same target centrality. These variations are valid and do not diminish performance; they simply reflect the natural diversity of approaches to a single problem. However, there were errors observed during the Analysis phase. For example, in Question 3, the Data Expert misinterpreted an average recovery count below one as a fraction rather than a discrete number of individuals. However, the Data Scientist detected this inconsistency and deferred to the Vision Expert, whose interpretation matched both the simulation protocol and analytical expectations. While such checks limit the impact of isolated calculation errors, occasional mistakes by DataExpert indicate that automated data analysis may be a part of further improvement.
@@
The framework also demonstrated sound awareness of its computational constraints. For example, in Question 3, we explicitly restricted simulations to the FastGEMF engine. The Simulation Scientist adhered to this limitation, employing FastGEMF exclusively for static networks, and it was explicitly noted in the final report that software limitations prevented temporal network experiments. This transparency is critical, as it prevents the framework from falsely claiming task completion. In contrast, such issues were observed in the single-agent implementation, which incorrectly claimed task accomplishment. Furthermore, the reliability of scientific outcomes often hinges on the algorithms employed, which can be inherently complex. For instance, in one of the trials in Question 3, the Simulation Scientist designed a custom engine to simulate disease spread over a temporal network but erred in updating the state ordering. This led to the premature recovery of initially infected nodes, preventing further spread and resulting in an underestimation of the final outbreak size and probability. Although the Data Scientist noted a significant discrepancy between simulated and expected analytical results, it incorrectly attributed the difference to the effects of temporality rather than the engine’s design flaw.
@@
To further investigate the agents’ attention to scientific details, in Question 3, we explicitly specified the constant activity rate (and not activity probability per unit time), $\alpha=3$, which is unlikely in real-world case studies and rare in literature, but was chosen to test the agents’ attention to subtle details. However, the Network Scientist incorrectly creates the network edges with the probability of a node being activated as $p=\alpha\Delta t$, with $\Delta t=1$, as the first order Taylor series approximation of $p=1-\exp^{-\alpha\Delta t}$. This approximation is acceptable only if $\alpha\Delta t\ll 1$ (However, if $\alpha$ is probability $\alpha\Delta t$ is exact). Therefore, it was invalid for $\alpha\Delta t=3>1$, resulting in all nodes being activated each time, instead of having exact probability $\approx 0.95$.
@@
1) the importance and impact of providing appropriate tools to obtain consistent and reliable results, as system’s lowest task success rate was in Question 3, which was not the most advanced but lacked both tools and expert knowledge and 2) the continued need for a human expert to validate results, because agents can still make mistakes that are difficult to detect and may produce persuasive responses that justify incorrect conclusions. In the next section, we categorize the system’s failure cases and error modes.
V-C1 Taxonomy of Failures and Errors
To rigorously analyze the gaps in the framework, we categorize failures into logic, assumption, algebra, data interpretation, and hallucination.
  • Logic or Methodological Error: Defined as unjustified logical leaps or invalid mathematical approximations, these errors were the most problematic as they often preserved workflow continuity while compromising scientific validity, and include about 60% of cases where the system failed to achieve the objective correctly. The stress-test study discussed in Question 3 was an example of an invalid approximation, resulting in a probability deviation of 0.05. Such a subtle logical error propagated undetected through the simulation and analysis phases, as the procedure remained syntactically correct. Another example is mapping errors, where severe logical mismatches were often detected. In one temporal network trial, the agent incorrectly mapped the analytical mean degree of a single snapshot ($k\approx 0.4$) to the aggregated static network model. This led to a sparse aggregated network and the isolation of 70% of nodes. Unlike the approximation error, this anomaly was diagnosed during the simulation phase through a reflection loop that prevented it from contaminating the final report. Algorithmic flaws are another form of logical error, such as premature recovery of infected individuals in the custom simulation engine used to spread over a temporal network. These errors are usually detected and resolved because they lead to unexpected results or execution errors, but they may still propagate silently, as in the aforementioned case, which resulted in a lower epidemic size.
  • Assumption Errors: These errors involve the introduction of convenient but incorrect premises that simplify the problem but invalidate the specific inquiry, and were about 30% of the task failure cases. For example, in targeted vaccination of Question 5, a failure mode was the assumption that the contact network followed a Poisson degree distribution in the discovery phase. While computationally convenient, this assumption caused the targeted vaccination of degree-10 nodes to be inefficient or impossible due to scarcity. In EpidemIQs, this failure was detected by designing two model setups and testing both the Poisson and a configuration model contact network. During the simulation and analysis phases, it was observed that the Poisson model yielded insufficient degree-10 nodes and correctly identified that the configuration model (with $\approx 11\%$ degree-10 nodes) was required to observe the threshold effect. In contrast, the iterative single-agent baseline consistently adhered to the Poisson assumption, which is a leading reason for the 0% task success rate on this objective.
  • Algebra and Data Interpretation Errors: While raw algebraic calculation errors were rare, we observed data interpretation errors during the analysis phase, usually due to complexity or large data sizes. For example, in the threshold vaccination question, the Data Expert occasionally misclassified vaccinated individuals as recovered, leading to a larger final size, mainly because in that SIR model setup, R included both initially vaccinated and recovered nodes. These errors were typically rare and were often corrected during ReAct refinement or cross-validation using Vision Expert.
  • Hallucination: Defined as confident output that is not true or not supported by evidence, was minimal in EpidemIQs’ results and strictly limited to time-to-time citation hallucination, generating plausible but non-existent references to support its claims. This stands in sharp contrast to the iterative single-agent baseline, which frequently exhibited process hallucination, particularly regarding task accomplishment when it had actually failed, such as claiming to have run a stochastic simulation when it had only solved a mean-field ODE.
Our classification aligns with recent evaluations of LLM reasoning capabilities [55], which identified logic and assumption errors as the primary obstacles in complex problem-solving tasks of LLMs. Our evaluation reveals that the hierarchical scientist-expert and multi-phase design successfully mitigated failure cases and demonstrated higher failure containment than iterative single-agent, but remains vulnerable to subtle errors, approximations, or assumptions that can propagate across phases.
V-D Limitations and Future Directions